### v2

# Plugging Attention into Power Grids: Towards Transparent Forecasting

Abstract
Reliable prediction of electricity demand plays a key role in safeguarding grid stability and guiding generation decisions, a need that grows with the decentralization and complexity of modern systems. While classical approaches such as Generalized Additive Models (GAMs) remain widely used, they often fail to capture the spatial dependencies inherent in energy networks. Graph Neural Networks (GNNs) offer a principled framework to incorporate this structure by directly leveraging graph topologies. In this work, we evaluate a broad set of GNN architectures —including GCN, GraphSAGE, ChebConv, TAG, APPNP, TransformerConv, and Graph Attention Networks (GAT and GATv2)— on two real-world electricity consumption datasets from France and the UK. Our results show that simpler models such as GCN, SAGE, or APPNP often outperform more complex alternatives in low-data regimes, while GAT ranks among the strongest architectures in our benchmarks, combining high accuracy with valuable interpretability. We perform a temporal analysis of attention weights, revealing evolving patterns of regional interaction linked to seasonal and meteorological variability. These results highlight that, although attention is not universally superior, it provides valuable explanatory power when spatial dependencies are prominent. Additionally, we demonstrate that ensemble-based expert aggregation strategies, particularly bottom-up combinations, significantly improve robustness and yield state-of-the-art performance across both datasets. These findings highlight the dual promise of GNNs for accurate and interpretable forecasting, and suggest that architectural simplicity coupled with ensemble methods can provide a practical path forward for transparent energy analytics.
Keywords:
Graph Neural Networks Attention Mechanisms Load Forecasting Expert Aggregation Interpretable Machine Learning
1 Introduction
The stability of power systems depends critically on maintaining an equilibrium between generation and demand, a task complicated by the fact that electricity cannot be stored efficiently and must therefore be produced in real time. While ongoing research explores ways to enhance the flexibility of various generation technologies, including nuclear energy [5], accurate short-term demand forecasts remain indispensable for market participants and system operators. The transition toward decentralized grid architectures introduces novel sources of uncertainty, adding complexity to forecasting tasks. At the same time, the growing share of intermittent renewable resources, such as wind and solar power, amplifies variability at multiple spatial scales due to the heterogeneous distribution of production sites. External shocks, notably the Covid-19 crisis and subsequent economic downturn, have further disrupted traditional consumption patterns, underscoring the importance of methods capable of handling non-stationary behaviors [2]. Fortunately, the increasing availability of geolocated and disaggregated consumption data provides new opportunities for models that can leverage such information to reduce forecast uncertainty [31, 40]. Parallel advances in adaptive forecasting techniques have also shown promise, demonstrating measurable improvements in predictive accuracy for aggregate electricity demand [7, 3].
Deep learning approaches have proven remarkably versatile in energy-related applications, showing strong capacity to adapt to diverse forecasting problems [29]. Achieving performance gains over traditional statistical models typically requires access to large datasets as well as the ability to capture complex relationships between predictors and targets. Such conditions are particularly present in decentralized power systems, where numerous forecasting tasks can be decomposed and modeled through suitable neural architectures. Among these, recurrent neural networks (RNNs), and especially Long Short-Term Memory (LSTM) networks, have been extensively applied to electricity demand. For instance, [34] introduced a deep LSTM framework for short-term household load forecasting, demonstrating substantial improvements compared with ARIMA models, while [28] proposed a sequence-to-sequence variant that successfully exploits the sequential structure of load time series. GNNs are increasingly recognized as effective architectures for time-series forecasting in energy systems, thanks to their ability to jointly represent spatial and temporal dependencies. Recent advances illustrate their versatility across different contexts. For instance, [36] proposed GLFN-TC, a short-term forecasting model that combines GNN layers with dilated temporal convolutions and residual connections, enabling automatic graph construction and improved predictive performance. In a multi-building setting, [30] designed a temporal GNN incorporating meteorological variables, which achieved competitive accuracy (MAPE of 3.2% at 15 minutes and 4.2% at 24 hours) while remaining robust to missing data. At a larger scale, PowerGNN [37] explicitly leverages network topology by combining GraphSAGE with GRUs, outperforming classical baselines in renewable-intensive scenarios. Beyond these domain-specific developments, earlier contributions in spatio-temporal learning laid the groundwork for applying GNNs to forecasting tasks. In traffic modeling, [16] introduced an attention-based architecture capable of dynamically weighting spatial and temporal information, while [10] proposed a model where the graph structure itself evolves over time to better capture dynamic dependencies. These ideas have since been transferred to the energy domain. For example, [23] developed a hybrid GNN for the Baidu KDD Cup 2022 wind power challenge, demonstrating the capacity of graph-based approaches to capture intricate inter-node relations in large-scale systems. More recently, [22] proposed an extended framework integrating contextual information, further enhancing the robustness and interpretability of load forecasting models. In earlier work, we investigated the use of GNNs for electricity consumption forecasting [9]. That study suggested that graph-based representations of aggregated load data can provide improvements over classical baselines, while also offering insights into spatial structures across regions. Attention mechanisms have become a central component in deep learning, particularly in sequence modeling and structured data processing. Initially introduced in the context of machine translation [4, 38], attention enables models to dynamically focus on the most relevant parts of the input. In the context of graph machine learning, Graph Attention Networks (GAT) [39] extend this idea by assigning learnable weights to neighboring nodes, allowing for adaptive neighborhood aggregation. More recent variants, such as GATv2 [8] and TransformerConv [35], aim to address limitations of the original formulation by improving expressiveness or robustness. Despite their popularity, several works have raised concerns about the interpretability of attention weights [21], highlighting the need for careful evaluation when using them as explanatory tools. In this work, we leverage attention scores not only for improving predictions but also for extracting interpretable spatio-temporal patterns in power networks. These contributions highlight the promise of GNNs in energy forecasting. However, a key challenge remains: designing or inferring graph structures that are tailored to the forecasting task, rather than relying on fixed or physically-defined topologies.
In this work, we make the following contributions:
  • We propose a graph-based forecasting framework for electricity consumption that integrates a broad spectrum of GNN architectures, including attention-based, diffusion-based, and spectral convolution models.
  • We observe that attention coefficients from GATs vary across time and exhibit structure when projected into low-dimensional spaces using PCA or UMAP. These projections reveal seasonal clusters that align with broad weather regimes (e.g. winter, summer). While these patterns do not constitute strict causal explanations, they offer a first step toward interpretable representations of temporal structure in graph-based forecasts and may support downstream tasks such as regime change detection or model analysis.
  • We conduct a systematic benchmark on two complementary datasets: a high-resolution UK dataset reflecting residential-level consumption, and a coarser French dataset with regional load measurements. We evaluate a wide array of GNNs, including GCN, GraphSAGE, GAT (v1/v2), TransformerConv, ChebConv, TAG, and APPNP.
  • We analyze the trade-offs between model complexity, interpretability, and predictive performance, showing that simpler models such as GCN or APPNP can outperform more complex ones in low-data regimes, while attention-based models like GAT remain competitive and uniquely valuable for explainability.
2 Preliminaries and Background
In this section, we introduce the basics of GNNs and the various convolutions used in the study.
2.1 Graphs and Message Passing
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph, where nodes $v\in\mathcal{V}$ represent entities and edges $(u,v)\in\mathcal{E}$ represent their interactions. The graph can be encoded via a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $A_{uv}=1$ indicates an edge between $u$ and $v$, or by a weighted matrix $\mathbf{W}$ allowing finer modeling of connection strengths. The neighborhood of node $v$ is denoted $\mathcal{N}_{v}=\{u\in\mathcal{V}~|~W_{uv}>0\}$. GNNs rely on message passing, a core mechanism where nodes iteratively aggregate and transform information from their neighbohood. At each layer $\ell$, the representation $\mathbf{h}_{v}^{(\ell)}$ of a node $v$ is updated based on its own previous embedding and those of its neighbors:
$\mathbf{h}_{v}^{(\ell)}$ $=\mathsf{UPDATE}^{(\ell)}\left(\mathbf{h}_{v}^{(\ell-1)},~\mathsf{AGGREGATE}^{(\ell)}\left(\left\{\mathbf{h}_{u}^{(\ell-1)}\,|\,u\in\mathcal{N}_{v}\right\}\right)\right),\ \mathbf{h}_{v}^{(0)}$ $=\mathbf{X}_{v},$ (1)
where AGGREGATE typically involves a permutation-invariant function (e.g. sum, mean, max) and UPDATE applies a nonlinear transformation. Stacking $d$ layers allows each node to access information from its $d$-hop neighborhood. This framework can be interpreted as a generalization of convolution to graph structures: message passing acts as a graph convolution operator [27], extracting local features while respecting the graph topology. This relational inductive bias enables GNNs to capture complex dependencies and structural patterns, making them well-suited for learning tasks on graph-structured data, such as node classification or forecasting over spatial networks [11].
2.2 Graph Convolutions
GCN.
The Graph Convolutional Network (GCN) [25] generalizes classical convolutions to non-Euclidean domains by aggregating feature information from local neighborhoods via the normalized adjacency matrix. This model is computationally efficient and provides a solid baseline for many graph learning tasks. However, its expressiveness is limited due to the fixed nature of the aggregation scheme and its inability to distinguish structural roles. The update rule is:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left(\sum_{u\in\mathcal{N}_{v}}\frac{1}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{v}|}}\mathbf{W}^{(\ell)}\mathbf{h}_{u}^{(\ell)}+\mathbf{b}\right).$
GraphSAGE.
GraphSAGE (SAmple and aggreGatE) [18] is an inductive GNN framework that learns a function to generate node embeddings by sampling and aggregating features from a node’s local neighborhood. Its ability to generalize to unseen nodes makes it especially suitable for dynamic or large graphs. The choice of aggregation function significantly impacts performance; we use the max-pooling aggregator in our experiments. Its update rule is:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left(\mathbf{W}^{(\ell)}\left[\mathbf{h}_{v}^{(\ell)}\bigg{|}\bigg{|}\max_{u\in\mathcal{N}_{v}}\left\{\sigma\left(\mathbf{W}_{\text{pool}}\mathbf{h}_{u}^{(\ell)}+\mathbf{b}\right)\right\}\right]\right),$
where $\left[\cdot||\cdot\right]$ is a concatenation operator.
GAT.
The Graph Attention Network (GAT) [39] incorporates an attention mechanism to assign different weights to different neighbors during message passing. This allows the model to learn the relative importance of neighboring nodes dynamically, offering better performance on heterogeneous graphs. However, GAT suffers from training instability and increased computational cost due to multi-head attention. The node update is defined as:
$\mathbf{h}_{v}^{(\ell+1)}=\bigg{|}\bigg{|}_{k=1}^{K}\sigma\left(\sum_{u\in\mathcal{N}_{v}}\alpha_{uv}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{u}^{(\ell)}+\textbf{b}\right),$
where the attention coefficients $\alpha_{uv}^{(k)}$ are computed via a softmax over neighbors based on a shared attention mechanism, and $K$ is the number of attention heads.
GATv2.
GATv2 [8] extends GAT by improving the expressiveness of the attention mechanism. It overcomes the static linear projection limitation of the original GAT by allowing the attention score to depend on both the source and target features after the linear transformation. This offers improved flexibility and results in many tasks. The node update remains similar to GAT, with the attention score computed as $\alpha_{uv}^{(k)}=\text{softmax}_{u}\left(\text{LeakyReLU}((\mathbf{a}^{(k)})^{\top}\mathbf{W}^{(k)}[\mathbf{h}_{v}\,||\,\mathbf{h}_{u}])\right).$
TransformerConv.
TransformerConv [35] adapts the Transformer architecture to graph data by using masked self-attention, restricting the attention mechanism to a node’s local neighborhood. This approach combines the benefits of global attention with graph topology, allowing richer contextual representations. While powerful, it requires careful regularization and is more computationally demanding. The update rule is:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left(\sum_{u\in\mathcal{N}_{v}}\text{softmax}_{u}\left(\frac{(\textbf{Q}\mathbf{h}_{v}^{(\ell)})^{\top}(\textbf{K}\mathbf{h}_{u}^{(\ell)})}{\sqrt{d}}\right)\textbf{V}\mathbf{h}_{u}^{(\ell)}+\textbf{b}\right),$
where Q, K, and V are learnable projection matrices.
TAGConv.
The Topology Adaptive Graph Convolutional Network (TAGConv) [13] uses a fixed-size filter constructed from powers of the adjacency matrix to capture multi-hop neighbor information. It provides a balance between spectral and spatial approaches and performs well on graphs where neighborhood depth is important. However, the fixed propagation depth may limit its flexibility. Its update rule is:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left(\sum_{k=0}^{K}\left(\textbf{D}^{-1/2}\textbf{A}\textbf{D}^{-1/2}\right)^{k}\mathbf{h}_{v}^{(\ell)}\mathbf{W}_{k}\right),$
where $K$ is the number of hops and $\mathbf{W}_{k}$ are learnable weights.
ChebConv.
ChebConv [12] approximates spectral convolutions using Chebyshev polynomials of the graph Laplacian. It allows for localized filters with adjustable receptive fields. Its main limitation lies in its dependency on spectral properties, which may not generalize well across different graph structures. The update rule is:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left(\sum_{k=0}^{K}T_{k}(\tilde{\textbf{L}})\mathbf{W}_{k}\right),$
where $T_{k}$ are Chebyshev polynomials and $\tilde{\textbf{L}}$ is the scaled Laplacian, with $T_{0}(\tilde{\textbf{L}})=\mathbf{h}_{v}^{(\ell)}$ and $T_{1}(\tilde{\textbf{L}})=\tilde{\textbf{L}}\mathbf{h}_{v}^{(\ell)}$.
APPNP.
The Approximate Personalized Propagation of Neural Predictions (APPNP) [15] decouples feature transformation from information propagation by applying a fixed personalized PageRank diffusion [32] to the output of a neural network. This design mitigates the risk of over-smoothing often encountered in deep message passing networks and improves the model’s stability, particularly in transductive and semi-supervised learning scenarios. After computing an initial node embedding via a multilayer perceptron $\mathbf{H}^{(0)}=\text{MLP}(\mathbf{X})$, the node features are propagated through the graph using the iterative update rule:
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left((1-\alpha)\sum_{u\in\mathcal{N}_{v}}\hat{A}_{uv}\mathbf{h}_{u}^{(\ell)}+\alpha\mathbf{h}_{v}^{(0)}+\mathbf{b}\right),$
where $\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ denotes the symmetrically normalized adjacency matrix with added self-loops, and $\alpha\in[0,1)$ is the teleport (or restart) probability that controls the strength of the personalized component in the propagation. In our application, we omit the MLP transformation and directly set $\mathbf{H}^{(0)}=\mathbf{X}$, i.e. the initial node features [^4] Pytorch Geometric/APPNP, which simplifies the architecture while preserving the core propagation mechanism of APPNP.
3 Graph Construction and Model Explainability
In this section, we outline our approach to building GNN-based forecasting models, emphasizing two key aspects: the construction of graph structures from different data sources and the interpretation of the resulting forecasts through explainability techniques.
3.1 Graph Construction from Data
Geographical similarity.
To capture the relationships between the nodes of a graph, a first approach is to study the similarity matrix of the geographical positions of these nodes. Let the kernelized geodesic distance between nodes $u$ and $v$ be $\mathbf{d}(u,v):=\exp\left\{-\frac{\mathbf{dist}(u,v)^{2}}{\sigma_{b}^{2}}\right\}$, where $\mathbf{dist}(u,v)$ is the geodesic distance and $\sigma_{b}$ is the kernel bandwidth. Then, we define the weight matrix as follows:
$\boldsymbol{W}_{\lambda}=(\boldsymbol{W}_{uv})_{u,v\in\mathcal{V}}=\mathbf{d}(u,v)\cdot\mathds{1}{\{\mathbf{d}(u,v)\geq\lambda\}}.$ (2)
In practice, $\lambda$ is picked such that the graph induced by $\boldsymbol{W}_{\lambda}$ is minimally connected, and $\sigma_{b}$ is set equal to the median pairwise geodesic distance.
Load and meteorological signals.
Rather than applying a dimensionality reduction technique, such as the singular value decomposition (SVD) used in [9], we retain only the most informative signal —the electricity load— to represent each region. This yields a global matrix of dimension $n\times T$ that captures the temporal dynamics of consumption across nodes and serves as a basis for subsequent graph construction algorithms. In this study, we evaluated three approaches for constructing the weight matrix based on the temporal similarity of regional load signals. The first method computes the Dynamic Time Warping (DTW) distance between regions using the FastDTW algorithm [33], which provides an efficient approximation with linear time and memory complexity. The second method relies on the empirical Pearson correlation matrix computed over the same time-series, where higher correlations suggest stronger connectivity. Finally, the third method uses the inverse of the covariance matrix —known as the precision matrix— which captures partial correlations and is particularly useful for identifying conditional dependencies between nodes. Each of these approaches yields a weighted adjacency matrix that can be used to define the graph structure for subsequent learning.
3.2 Explainability
In [9], the GNNExplainer algorithm has been used for the identification of compact subgraphs that best explain the model’s predictions, by maximizing the mutual information between the explanatory subgraph and the forecast. In this work, we adopt a complementary approach based on the analysis of attention weights over time. These weights, defined for each pair of nodes, provide an interpretable proxy for the flow of information across the graph. Considering the questions put by prior work on the reliability of attention weights as faithful explanations [21], we do not treat them as definitive indicators of causal influence. Instead, we study the distribution of attention matrices projected into lower-dimensional spaces (via PCA and UMAP) under varying meteorological conditions. This approach enables us to reveal recurrent patterns of regional similarity and detect shifts in interaction structures across seasons.
4 Experiments
In this section, we outline the experimental procedure used to assess GNNs, covering everything from dataset preparation to result analysis, including model parameterization.
4.1 Datasets
French Dataset.
We evaluate GNN models on regional electricity consumption data provided by RTE (Réseau de Transport d’Électricité), the French electricity transmission system operator. The dataset spans from January 2015 to December 2019 and includes various features relevant to electricity demand and meteorological conditions across French administrative regions. Calendar-related features include a trend index, multiple date encodings, month and year indicators, time-of-day and day-of-year information, week numbers, day type (weekday, weekend, public holiday), specific holiday periods (e.g. Christmas, summer, national holidays), and a daylight saving time indicator. Meteorological features (e.g. temperature, wind, solar irradiance) are averaged across the 32 national weather stations and spatially aggregated by administrative region, providing region-level summaries of environmental conditions. The target variable is the national electricity load (in megawatts). We also include lagged consumption features, as detailed in Sec. 4.2.
UK Dataset.
The UK dataset [^5] https://weave.energy/ consists of aggregated half-hourly residential electricity consumption data collected via smart meters by four UK Distribution Network Operators (DNOs). Currently, data are available from the two largest DNOs, Scottish and Southern Electricity Networks (SSEN) and National Grid Electricity Distribution (NGED), covering approximately 2 million smart meters, $120,000$ low voltage feeders, and $50,000$ substations, starting from January 2024 (NGED) and February 2024 (SSEN), with updates subject to publication delays of 5 to 30 days. The data primarily cover regions such as Northern Scotland, the Midlands, South Wales, South West England, and Southern England. In our experiments, we use only calendar-based features (e.g. day of the week, time of day) and lagged consumption values, consistent with the setup used for the French dataset. The target variable corresponds to the electricity demand measured at residential substations.
Synthetic Dataset.
In order to analyze the behavior and interpretability of attention weights in GATs, we generate synthetic datasets specifically designed for structural analysis rather than forecasting performance. Each dataset consists of $N$ time-series, one per node in the graph, defined as the sum of a smooth trend, a periodic component, and additive high-frequency noise. The noise is intentionally strong to prevent models from simply fitting the individual signals, thereby encouraging the discovery of meaningful cross-node dependencies. To simulate interactions between nodes, we transform the original input signals $\mathbf{X}\in\mathbb{R}^{N\times T}$ by $\mathbf{A}\mathbf{X}$, where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a coupling matrix. The supervised learning objective is to reconstruct these signals from the initial inputs. This ensures that the predictive performance essentially hinges on leveraging the coupling structure implicitly embedded in the data. We construct three variations of synthetic data:
  • In the single coupling setting, a fixed, sparse, full-rank matrix $\mathbf{A}$ is used at every time step, and no additional input is provided.
  • In the explicit switching setting, two distinct coupling matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ are used at different time steps, with an binary input indicating the active matrix ($1$ for $\mathbf{A}_{1}$, 0 for $\mathbf{A}_{2}$). The switching signal is generated as a binary random vector of length $T$, where each value is independently drawn from a uniform distribution over $\{0,1\}$. This results in a stochastic alternation between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ at each time step (e.g. 0101110101001).
  • In the ambiguous switching setting, the same alternation between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used, but the model is now provided with two input categorical variables that only jointly (hence indirectly) encode which coupling matrix is in effect each time. Specifically, under $\mathbf{A}_{1}$, the pair takes a value of either $(0,0)$ or $(1,1)$ at random, while under $\mathbf{A}_{2}$, the pair takes a value of either $(0,1)$ or $(1,0)$ at random. This design ensures that each individual variable is uninformative on its own, and only their joint value contains the relevant information about the active coupling.This setup creates ambiguity, mimicking real-world latent factors that are not directly observable.
4.2 Experimental Settings
We consider a multivariate regression setting in which the objective is to forecast the electricity consumption for each node over the course of the following day, based on the features observed during the current day. The model takes as input a sequence of feature vectors covering the past $48$ half-hour intervals and predicts the consumption values for the next $48$ intervals that corresponds to the next day’s demand profile.
Training Procedure.
Let $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times d\times T}$ denote the tensor of normalized historical node features, where $n$ is the number of nodes, $d$ the number of features per node, and $T$ the number of time steps. The target values are represented by $\mathbf{y}\in\mathbb{R}^{n\times T}$. The forecasting problem is cast as a sequence-to-sequence regression task with an input window of length $w=48$ and an output horizon of length $h=48$. At each time step $t$, the GNN $\Phi_{\boldsymbol{\theta}}$ defined on the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ maps the current covariates $\tilde{\mathbf{X}}_{t}$ together with the recent target history $\mathbf{y}_{t-w:t-1}$ to predictions $\hat{\mathbf{y}}_{t:t+h-1}=\Phi_{\boldsymbol{\theta}}(\mathcal{G},\tilde{\mathbf{X}}_{t}\cup\mathbf{y}_{t-w:t-1})$. Model parameters are trained by minimizing the mean squared error (MSE) between predicted and ground-truth sequences, i.e. $\boldsymbol{\theta}^{*}\in\arg\max_{\boldsymbol{\theta}}\;\sum_{t\in\mathcal{T}}\lVert\Phi_{\boldsymbol{\theta}}(\mathcal{G},~\tilde{\mathbf{X}}_{t}\cup\mathbf{y}_{t-1:t-w})-\mathbf{y}_{t:t+h-1}\rVert^{2}$. The parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer [24], which provides adaptive learning rates and momentum-based updates for improved convergence. We adopt distinct data splits for the two datasets: for the French dataset, the training period spans from January 1, 2015 to December 31, 2017, followed by a one-year validation set (2018) and a one-year test set (2019). For the UK dataset, we use a shorter window: 10 days of training (February 13 to February 22, 2024), one day for validation (February 23, 2024), and three days for testing (February 24 to 26, 2024).
Baseline models.
We evaluated three baseline forecasting strategies to predict the aggregated electricity demand at multiple consumption nodes. The first approach trains a separate feedforward neural network for each node to forecast the next 48 half-hourly values (i.e. the next day), using all available information up to time t1t-1. The second baseline is a classical statistical method: we fit 48 independent SARIMA models per node— one for each half-hour slot —using only historical values (lags), without any exogenous covariates. This approach follows standard practices in time-series modeling that emphasize seasonality and decomposition [6]. The third baseline is a simple yet competitive persistence method, where the forecast for each time slot is directly copied from the same half-hour either on the previous day (D1D-1) or the previous week (D7D-7), a common benchmark in load forecasting studies [19]. These baselines provide a diverse set of references spanning machine learning, statistical, and naive forecasting paradigms. While GAMs are widely used in electricity forecasting for their ability to model nonlinear effects of multiple covariates [9], we did not include them here due to the limited availability of informative exogenous features in our setting —particularly in the UK dataset, where only calendar features and lagged consumption values are available. In such contexts, the strength of GAMs in capturing smooth covariate effects becomes less relevant, and simpler autoregressive strategies often suffice [17].
Evaluation Metrics.
We evaluate forecasting performance at the top scale using two standard loss functions: Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE). MAPE offers an interpretable percentage-based error metric, useful for comparing across scales:
$\mathsf{MAPE}(\mathbf{y},\hat{}\mathbf{y})=\dfrac{1}{T}\sum_{t=1}^{T}\left|{\textstyle\dfrac{\sum_{i=1}^{N}y_{i,t}-\hat{y}_{i,t}}{\sum_{i=1}^{N}y_{i,t}}}\right|\qquad\qquad\mathsf{RMSE}(\mathbf{y},\hat{}\mathbf{y})=\sqrt{\dfrac{1}{T}\sum_{t=1}^{T}\left(\sum_{i=1}^{N}y_{i,t}-\hat{y}_{i,t}\right)^{2}}$ (3)
RMSE penalizes large deviations and is suited for assessing critical errors in load prediction [20].
Parametrization.
To ensure fair and performant comparisons across the studied graph-based architectures, we optimized the model hyperparameters using the Optuna library [1]. The search space included core architectural parameters common to all models, such as the number of layers, the number of hidden channels per layer, the batch size, and the learning rate. Each of these plays a distinct role: n_layers controls the receptive field and model depth, hidden_channels regulates the model capacity, batch_size affects the training stability and generalization, and lr determines the speed and stability of convergence. We further included model-specific parameters to adapt the search space to each architecture. For spectral and spatial methods, such as ChebConv and TAGConv, we optimized the neighborhood order (e.g. Chebyshev polynomial degree or number of hops), which modulates the locality of the convolutional filters. For attention-based models such as GAT and TransformerConv, we additionally tuned the number of attention heads, which controls the diversity of subspaces in which attention scores are computed. For APPNP, we included the number of propagation steps $K$, which determines the extent of information diffusion, and the teleport (or restart) probability $\alpha$, which balances the trade-off between local propagation and retention of the original signal. All experiments were performed using early stopping based on validation loss, and each configuration was evaluated through multiple runs. This principled tuning procedure enabled the selection of performant and robust architectures across both datasets. It is also important to note that the graph construction methods differ across datasets: in the French dataset, regional aggregation allows for the meaningful use of spatial graphs informed by physical proximity or meteorological similarity. In contrast, for the UK dataset —which involves highly disaggregated residential-level signals— using a fixed spatial graph can be counterproductive, as the spatial structure is weakly informative or even noisy. In this case, we relied on data-driven graph construction strategies better suited to fine-grained forecasting. The results of the optimization are reported in Table 1.
Table 1: Hyperparameters of the GNN models for the French and the UK dataset.
Dataset Model Graph structure batch_size n_layers hidden_channels lr Other HP
[…]
SAGE Correlation 16 2 401 $4\cdot 10^{-4}$ None
GAT GL3SR 16 3 364 $1\cdot 10^{-3}$ $\texttt{heads}=2$
GATv2 Space 16 2 359 $4\cdot 10^{-4}$ $\texttt{heads}=2$
Transformer DistSplines 16 1 335 $3\cdot 10^{-3}$ $\texttt{heads}=2$
TAG DTW 16 3 329 $3\cdot 10^{-4}$ $K=1$
Cheby DistSplines 16 1 249 $6\cdot 10^{-4}$ $K=3$
APPNP Space 16 1 106 $2\cdot 10^{-3}$ $(K,\alpha)=(1,0.93)$
[…]
SAGE Correlation 32 5 207 $7\cdot 10^{-4}$ None
GAT Correlation 16 1 172 $3\cdot 10^{-3}$ $\texttt{heads}=1$
GATv2 Precision 128 3 71 $2\cdot 10^{-3}$ $\texttt{heads}=1$
Transformer Correlation 32 5 354 $4\cdot 10^{-4}$ $\texttt{heads}=1$
TAG DTW 64 1 70 $3\cdot 10^{-3}$ $K=4$
Cheby Correlation 32 4 118 $6\cdot 10^{-2}$ $K=10$
APPNP Precision 64 3 454 $2\cdot 10^{-2}$ $(K,\alpha)=(8,0.85)$
Uncertainty Quantification and Model Aggregation.
A simple yet effective way to quantify uncertainty in neural networks is to train multiple models independently with different random initializations and aggregate their predictions [26]. This approach, known as Deep Ensembles, captures epistemic uncertainty —i.e. uncertainty stemming from limited knowledge or data— by leveraging the variability in predictions induced by different parameter configurations. It is particularly well-suited when models exhibit high variance, than can happen if small amounts of data are available or the variables are highly correlated. Given that each forecasting model carries distinct strengths and inductive biases, we further exploit their complementarity through expert aggregation. In particular, we employ the ML-Poly algorithm [14], a robust online learning method that dynamically adjusts expert weights based on past forecasting performance. In our setting, aggregation is performed every $48$ time steps (i.e. once per daily prediction horizon), with weights updated daily to reflect recent accuracy. Formally, for time step $t$, let $x_{j,t}$ denote the forecast of expert $j$ and $p_{j,t}$ its weight; the aggregated prediction is given by: $\widehat{y}_{t}=\sum_{j=1}^{K}p_{j,t}x_{j,t}$ where $K$ is the number of experts. To avoid overfitting and reduce the complexity of model selection, we also consider uniform aggregation, which consists of a simple average over all expert forecasts. Despite its simplicity, this baseline often performs competitively —sometimes even outperforming the best individual expert— thanks to the regularizing effect of ensemble averaging which reduces the variance of predictions.
4.3 Results
Numerical Results.
On the French dataset, which has a coarser spatial resolution and includes more aggregated load profiles at the regional level, performance across GNN architectures is relatively homogeneous. With uniform aggregation (over 10 random initializations for each method), SAGE delivers the best RMSE ($873$ MW) and a competitive MAPE ($1.12\%$), while the vanilla GAT achieves the lowest MAPE ($1.11\%$), confirming its ability to balance expressiveness and stability. Simpler models such as GCN and TAG remain highly competitive, and the gains from more complex architectures like GATv2 and TransformerConv are marginal or even negative. These findings suggest that in moderately sized graphs with stable and aggregated targets, moderately expressive models can match or outperform more elaborate architectures, particularly when using uniform aggregation, which helps mitigate overfitting risks and reduces variance across experts. Baseline methods such as FF, SARIMA, or persistence approaches are consistently outperformed by GNN-based models across both datasets, particularly in terms of RMSE. This highlights the benefit of exploiting graph structure for load forecasting, even though simpler baselines remain competitive in some cases (e.g., FF on the French dataset, and SARIMA on the UK dataset).
In contrast, the UK dataset is characterized by a much finer granularity and a greater number of disaggregated nodes (residential substations), where learning is more sensitive to data sparsity and noise. Under these conditions, APPNP with uniform aggregation achieves the best results, reaching the lowest MAPE ($6.54\%$) and RMSE ($12.61$ MW), followed by GAT ($7.79\%$, $14.13$ MW) and SAGE ($8.19\%$, $14.87$ MW). In this data-scarce regime, complex attention-based architectures such as GATv2 and TransformerConv perform worse, likely due to their higher parameterization and the resulting overfitting. The strong performance of APPNP can be attributed to its propagation scheme with teleportation, which regularizes diffusion and stabilizes learning. GATs’ strong results confirm that attention mechanisms can still offer competitive accuracy, especially when combined with uniform averaging, which smooths predictions across models and limits overfitting. These observations reinforce the value of simple and stable GNN designs in fine-grained and noisy environments.
Overall, the results highlight the importance of architectural simplicity and strong inductive biases in low-data, high-resolution scenarios, as well as the value of graph-aware models over global baselines. Across both datasets, the choice of aggregation strategy significantly influences performance, often surpassing the differences between GNN architectures themselves. On the French dataset, the bottom aggregation consistently yields the best results, achieving the lowest MAPE ($1.01\%$) and RMSE ($789$ MW). This suggests that learning to combine expert predictions at the local (node) level before aggregating them captures more informative patterns than naive averaging. Similarly, on the UK dataset, where signals are more heterogeneous and data is sparser, bottom aggregation also outperforms other strategies with the lowest MAPE, with $7.68\%$, and RMSE, with $14.35$ MW. This supports the idea that local aggregation provides a finer control over uncertainty and variation at the node level before summarizing across the network. In contrast, top aggregation —which combines predictions after summation— tends to perform worse, especially in the UK setting, likely because it neglects node-level diversity. Moreover, while uniform weighting remains competitive, particularly in simpler models, learned aggregation offers systematic gains, underlining the benefit of adaptive combination rules in ensemble-based forecasting.
Table 2: Average and best test performance per model – ‘Unif.’ indicates uniform averaging of expert predictions, ‘Agg.’ refers to online aggregation using the MLPol algorithm. ‘Bottom’ aggregation aggregates at the node level before summation, while ‘Top’ aggregation first sums expert predictions and then aggregates. ‘*’ means the models were only ran once as as their training procedure is deterministic.
Model French Dataset UK Dataset
MAPE (%) RMSE (MW) MAPE (%) RMSE (MW)
Avg. $\pm$ Std Best Avg. $\pm$ Std Best Avg. $\pm$ Std Best Avg. $\pm$ Std Best
GCN $1.75\pm 0.51$ $1.21$ $1255\pm 359$ $903$ $9.52\pm 1.39$ $6.91$ $16.68\pm 2.27$ $12.13$
Unif. $1.21$ $895$ $8.63$ $15.09$
Agg. $\underline{1.09}$ $\mathbf{839}$ $8.51$ $14.91$
SAGE $30\pm 0.09 1.1$41.14 $998\pm 72$ $862$ $8.63\pm 0.84$ $6.95$ $15.61\pm 1.49$ $13.10$
Unif. $\underline{1.12}$ $\mathbf{873}$ $8.19$ $14.87$
Agg. $1.14$ $898$ $\underline{8.07}$ $14.85$
GAT $42\pm 0.23 1.1$31.13 $1052\pm 136$ $888$ $10.01\pm 1.80$ $6.59$ $18.11\pm 2.96$ $12.79$
Unif. $\mathbf{1.11}$ $883$ $\underline{7.79}$ $\underline{14.13}$
Agg. $\mathbf{1.08}$ $\underline{871}$ $8.29$ $\underline{14.72}$
GATv2 $1.64\pm 0.34$ $1.22$ $1165\pm 200$ $937$ $11.32\pm 1.48$ $8.21$ $20.22\pm 2.53$ $15.23$
Unif. $1.14$ $888$ $8.78$ $16.44$
Agg. $1.14$ $902$ $8.94$ $16.40$
Transformer $1.50\pm 0.17$ $1.28$ $1123\pm 118$ $950$ $11.93\pm 1.41$ $9.69$ $21.24\pm 2.41$ $16.61$
Unif. $1.14$ $883$ $8.36$ $15.32$
Agg. $1.15$ $910$ $8.48$ $15.53$
TAG $1.36\pm 0.17$ $1.22$ $1022\pm 107$ $925$ $10.63\pm 1.16$ $7.69$ $18.35\pm 1.81$ $14.21$
Unif. $1.14$ $884$ $9.23$ $15.79$
Agg. $1.14$ $881$ $9.41$ $16.19$
Cheby $1.39\pm 0.14$ $1.23$ $1036\pm 91$ $925$ $9.49\pm 1.12$ $7.95$ $18.38\pm 1.83$ $15.66$
Unif. $1.16$ $\underline{878}$ $8.60$ $17.23$
Agg. $1.15$ $876$ $8.93$ $17.37$
APPNP $1.48\pm 0.17$ $1.23$ $1091\pm 114$ $931$ $8.31\pm 1.53$ $6.01$ $5.41\pm 2.73 11$.0111.01
Unif. $1.22$ $944$ $\mathbf{6.54}$ $\mathbf{12.61}$
Agg. $1.20$ $943$ $\mathbf{7.60}$ $\mathbf{14.13}$
$F$ $1.50\pm 0.06$ $1.41$ $1101\pm 31$ $1044$ $10.71\pm 0.40$ $9.83$ $18.48\pm 609$ $17.06$
SARIMA $^{*} 3.293.29$ $^{*} 26532653$ $^{*} 8.068.06$ $^{*} 16.6016.60$
Persistence–11 $^{*} 5.795.79$ $^{*} 45124512$ $^{*} 8.868.86$ $^{*} 17.2017.20$
Persistence–77 $^{*} 6.036.03$ $^{*} 46564656$ $^{*} 10.1910.19$ $^{*} 19.1919.19$
Uniform $1.06$ $836$ $8.10$ $14.74$
Bottom Aggregation $\mathbf{1.01}$ $\mathbf{789}$ $\mathbf{7.68}$ $\mathbf{14.35}$
Top Aggregation $1.02$ $817$ $8.33$ $15.04$
Figure 1: Violin plot of the RMSE error values on the French dataset for the GNN models.
Figure 2: Violin plot of the RMSE error values on the UK dataset for the GNN models.
Explainability Results.
To gain interpretability from attention-based architectures, we analyze the attention weights learned by GATs. For each time step in the dataset, the model computes a full attention matrix over all pairs of regions, quantifying the relative importance of inter-nodal interactions for the prediction task. These matrices, dynamically updated across time, encode rich spatio-temporal patterns. For instance, two regions may exhibit similar attention profiles in spring and autumn due to analogous consumption behaviors, yet diverge significantly in summer or winter —particularly in regions affected by seasonal tourism. To explore the structure embedded in these matrices, we apply dimensionality reduction techniques such as Principal Component Analysis (PCA) and Uniform Manifold Approximation and Projection (UMAP). These methods are especially useful for uncovering latent patterns in high-dimensional data: PCA captures directions of maximal variance, offering a linear projection, while UMAP preserves local and global topological structures, revealing non-linear relationships. As illustrated in Fig. 3, both projections reveal a clear seasonal organization of the attention matrices. The color gradient, indicating the day of the year, reveals well-separated clusters corresponding to different periods (e.g. winter, summer). This seasonal structuring suggests that the attention mechanism implicitly captures underlying physical variables such as temperature, daylight variation, or regional activity cycles —providing an interpretable signal beyond raw consumption values. These findings reinforce the potential of attention-based models not only as powerful predictors, but also as tools for understanding spatio-temporal dynamics in energy consumption.
(a) PCA.
(b) UMAP.
Figure 3: PCA and UMAP projections of attention matrices on the French dataset reveal seasonal clustering: spring (green), summer (blue), and winter (red).
(a) Explicit coupling: Layer 22.
(b) Implicit coupling: Layer 0.
(c) Implicit coupling: Layer 11.
(d) Implicit coupling: Layer 22.
Figure 4: PCA of attention matrices on the synthetic dataset with two coupling modes ($\mathbf{A}_{1}$ or $\mathbf{A}_{2}$). Colors reflect the coupling matrix in effect. (a) Explicit coupling information: the mode is explicitly given to the model as a binary input. (b—d) Implicit coupling information: the mode in effect is indicated implicitly by the combined values of two input binary variables. The subfigures show the PCA of attention matrices across GAT layers.
Interpreting Attention on Synthetic Data.
To better understand the nature of the information captured by attention mechanisms in GATs, we conduct a series of experiments using the synthetic datasets previously described. These datasets offer controlled conditions to interpret how attention matrices reflect inter-node interactions and latent structural cues. In the single coupling scenario, where a fixed matrix $\mathbf{A}$ governs all interactions, we observe that the average attention matrix learned by the model does not resemble the ground truth coupling matrix. Moreover, no obvious transformation seems to relate the two. This suggests that attention weights do not directly encode interpretable physical dependencies. Nevertheless, the low standard deviation of attention values across time indicates that the model learns a temporally stable pattern of interactions. To explore whether GATs can adapt to multiple interaction regimes, we consider a second scenario involving a switch between two coupling matrices, $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$. In the explicit switching scenario, the model is given as input a binary vector indicating which of those matrices is in effect at each moment. After training, PCA is applied to the attention vectors to reveal two clearly separable clusters, as shown in Fig. 4(a), demonstrating that the GAT can modulate its internal attention representations according to the active coupling mode.
To evaluate the model’s ability to recover hidden structure from ambiguous input, we design a third scenario with ambiguous switching. In this case, the same switching pattern between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used as before, but the information about which of those matrices is in effect each time is now implicitly encodes by two categorical variables. This setup introduces ambiguity by suggesting four categories where only two actual regimes exist. We perform a PCA of the learned attention vectors at each GAT layer. As shown in Fig. 4, the first layer exhibits no clear separation. In the second layer, three intermediate clusters emerge, corresponding to the four ambiguous exogenous categories. In the final layer, these are merged into two well-separated clusters, correctly matching the underlying coupling structure. This progression suggests that GATs can disentangle latent generative factors through hierarchical attention refinement, even in the presence of misleading categorical inputs. This property is particularly valuable in real-world forecasting scenarios, where exogenous variables may be noisy, redundant, or poorly specified —as is often the case with categorical features in energy systems. In such cases, the model’s ability to infer underlying structure without explicit supervision can enhance robustness and generalization. Moreover, it opens promising avenues for model interpretability: analyzing how attention distributions evolve across layers may help identify meaningful patterns or changes of regime, even when external signals are ambiguous or incomplete. Ultimately, this illustrates how attention-based GNNs can support both prediction and explanation, by learning to filter out misleading information and emphasizing latent, structure-relevant features.‘
5 Conclusion
This paper investigates the use of attention-based graph neural networks for electricity consumption forecasting, with a strong focus on interpretability. We show that attention mechanisms not only improve forecasting accuracy on certain datasets, but also provide insights into time-varying regional dependencies. Our analysis of attention weight distributions demonstrates that GAT-based architectures can partially recover meaningful seasonal and meteorological patterns from the data, and highlights how attention mechanisms provide a structured and interpretable lens through which temporal and spatial dependencies can be explored. Interestingly, our results also show that simpler architectures such as GCN and APPNP often outperform deeper or more complex models (e.g. GATv2, TransformerConv), especially in highly disaggregated, low-data settings such as the UK residential network. These findings emphasize the importance of model simplicity and strong inductive biases in real-world forecasting scenarios. Finally, we show that ensemble techniques —particularly bottom-up aggregations— offer a practical and effective strategy to improve robustness.
Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.